OpenStudy (anonymous):

solve: (step by step if you would) s^3 + 4s^2 - s - 4 = 0

OpenStudy (anonymous):

HINT: one of your solutions is s = 1. From there you can figure out the solution to the resulting quadratic.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

You'll need to factor that left side of the equation. Now the hint given was that s=1 is a root which means that (s - 1) is a factor of s^3 + 4s^2 - s - 4. If you take \(\large \frac{s^3+4s^2-s-4}{s-1} \), you'll get a quadratic factor which is also factorable.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i still need help.....

jhonyy9 (jhonyy9):

s^2(s +4) -(s+4) =(s^2 -1)(s+4)